OpenStudy (anonymous):

Simplify the rational expression. State any restrictions on the variable. \[\frac{n^{4}-10n^{2}+24}{n^{4}-9n^{2}+18}\]

OpenStudy (anonymous):

i guess the idea is to factor and cancel

OpenStudy (anonymous):

\[x^4-10x+24=(x^2-6)(x^2-4)\] for the numerator

OpenStudy (anonymous):

\[x^4-9x^2+18=(x^2-3)(x^2-6)\] for the denominator

OpenStudy (anonymous):

looks like you can cancel something, right?

OpenStudy (anonymous):

as for the "restrictions" set \[(x^2-6)(x^2-4)=0\] and solve for \(x\) since you are not allowed to divide by zero

OpenStudy (anonymous):

is x = 0 ? I'm not too familiar with solving for x to get restrictions...

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[(x^2-6)(x^2-4)=0\] \[x^2-6=0\\ x^2-4=0\] \[x^2-6=0\iff x^2=6\iff x=\pm\sqrt6\] \[x^2-4=0\iff x^2=4\iff x=\pm2\]

OpenStudy (anonymous):

I'm kinda confused about the second one because there's no two's in any othe answer choices http://awesomescreenshot.com/0772ddk722 unless... was I supposed to plug those into x?